A company has learned that by pricing a newly released toy at ​$7​, sales will reach 3000 per day. Raising the price to ​$9 will cause the sales to fall to 2000 per day. Assume that the ratio of change in price to change in daily sales is constant and let x be the price of the toy and y be the number of sales.

a. Find the linear equation that models the​ price-sales relationship for this toy.​ [Hint: The line must pass through ​(7​,3000​) and ​(9​,2000​).]

b. Use this equation to predict the daily sales of the toy if the price is set at ​$8.50.


1
Expert's answer
2022-02-17T18:34:56-0500

a.


"slope=m=\dfrac{y_2-y_1}{x_2-x_1}=\dfrac{2000-3000}{9-7}=-500"

"y=mx+b"

Substitute


"3000=-500(7)+b=>b=6500"

The linear equation that models the​ price-sales relationship for this toy is


"y=-500x+6500"

b. Given "x=\$8.50"


"y=-500(8.5)+6500"

"y=2250"

The daily sales of the toy will be 2250 per day.